OpenStudy (anonymous):

Solve for u. (-3/4u+12)+2=(-3/u+3)

ganeshie8 (ganeshie8):

start with multiplying 4u through out

OpenStudy (anonymous):

Is the problem \[-\frac{3}{4u + 12} + 2 = -\frac{3}{x + 3}?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh wait no. there is no x in the problem

OpenStudy (anonymous):

\[-\frac{3}{4u + 12} + 2 = -\frac{3}{u + 3}\]\[-\frac{3}{4(u + 3)} + 2 = -\frac{3}{u + 3}\]\[-\frac{3}{4(u + 3)} + (2 \times \frac{4(u + 3)}{4(u + 3)}) = (-\frac{3}{u + 3} \times \frac 44)\]Can you solve from here?

OpenStudy (anonymous):

no i still need help

OpenStudy (anonymous):

\[-\frac{3}{4(u + 3)} + (2 \times \frac{4(u + 3)}{4(u + 3)}) = (-\frac{3}{u + 3} \times \frac 44)\]\[-\frac{3}{4(u + 3)} + \frac{8(u + 3)}{4(u + 3)} = -\frac{12}{4(u + 3)}\]Drop the denominators.\[-3 + 8(u + 3) = -12\]Now can you solve?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

np :)